OpenStudy (anonymous):

Write the following statement as an equation and solve. Jake Drewrey, owner of Jake's Landscape Service, knows that one of his fertilizer tanks holds four times as many gallons of liquid fertilizer as a second tank. The two tanks together hold 371 gal. If both tanks are filled to capacity, how many gallons of fertilizer does each tank hold?

OpenStudy (anonymous):

i think.... its one tank holds 92.75 gallons another tank holds278.25 gallons... again i think im not for sure

OpenStudy (anonymous):

btw \(\cal\Huge\color{gold}{WELCOME~TO~OPENSTUDY!!!}\)

OpenStudy (anonymous):

hey diamond bear im sorry i left ur question my power went out for a while

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

thnx here

OpenStudy (anonymous):

do u still wanna learn latex?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

alright ill tag u

OpenStudy (anonymous):

sweet

OpenStudy (texaschic101):

x = 1st tank y = 2nd tank x + y = 371 x = 4y 4y + y = 371 5y = 371 y = 371/5 y = 74.2 x + y = 371 x + 74.2 = 371 x = 371 - 74.2 x = 296.8 1st tank (x) holds 296.8 gallons and 2nd tank (y) holds 74.2 gallons